### v1

# Sparse Attention Post-Training for Mechanistic Interpretability

Abstract
We introduce a simple post-training method that makes transformer attention sparse without sacrificing performance. Applying a flexible sparsity regularisation under a constrained-loss objective, we show on models up to 1B parameters that it is possible to retain the original pretraining loss while reducing attention connectivity to $\approx 0.3\%$ of its edges. Unlike sparse-attention methods designed for computational efficiency, our approach leverages sparsity as a structural prior: it preserves capability while exposing a more organized and interpretable connectivity pattern. We find that this local sparsity cascades into global circuit simplification: task-specific circuits involve far fewer components (attention heads and MLPs) with up to 100× fewer edges connecting them. These results demonstrate that transformer attention can be made orders of magnitude sparser, suggesting that much of its computation is redundant and that sparsity may serve as a guiding principle for more structured and interpretable models.
1 Introduction
Scaling has driven major advances in artificial intelligence, with ever-larger models trained on internet-scale datasets achieving remarkable capabilities across domains. Large language models (LLMs) now underpin applications from text generation to question answering, yet their increasing complexity renders their internal mechanisms largely opaque [Bommasani2021FoundationModels]. Methods of mechanistic interpretability have been developed to address this gap by reverse-engineering neural networks to uncover how internal components implement specific computations and behaviors. Recent advances in this area have successfully identified interpretable circuits, features, and algorithms within LLMs [nanda2023progress, olsson2022context], showing that large complex models can, in part, be understood mechanistically, opening avenues for improving transparency, reliability, and alignment [bereska2024mechanistic].
However, interpretability is bottlenecked by the model itself: even with sophisticated reverse-engineering techniques that can faithfully reveal internal algorithms, the underlying computations implemented by large models can still remain highly complex and uninterpretable. Circuits for seemingly simple tasks may span hundreds of interacting attention heads and MLPs with densely intertwined contributions across layers [conmy2023towards], and features can influence each other along combinatorially many attention-mediated paths, complicating attention attribution [kamath2025tracing]. To exemplify this, Figure 1 (top) illustrates the attention patterns of a small, single-head transformer trained solely on a simple two-digit addition task. Here, the model has learned to solve the task in a highly diffused manner, where information about each token is dispersed across all token locations, rendering the interpretation of the underlying algorithm extremely difficult even in this simple case.
The crux of the problem is that models are not incentivised to employ simple algorithms during training. In the present work, we advocate for directly embedding interpretability constraints into model design in a way that induces simple circuits while preserving performance. We focus our analysis on attention mechanisms and investigate sparsity regularisation on attention patterns, originally proposed in [lei2025spartan], as an inductive bias. To demonstrate how sparse attention patterns can give rise to interpretable circuits, we return to the two-digit addition example: Figure 1 (bottom) shows the attention patterns induced by penalizing attention edges during training. Here, the sparsity inductive bias forces the model to solve the problem with much smaller, intrinsically interpretable computation circuits.
Note that we apply this sparsity regularisation scheme as a post-training strategy for pre-trained LLMs. Because the method operates entirely post-training, it can be applied to existing models without re-running pretraining, offering a flexible way to sparsify their attention patterns and improving their interpretabiliy. We show, on models of up to 1B parameters, that the proposed procedure preserves the original cross-entropy loss on pretraining data while reducing the effective attention map to $\approx 0.3\%$ of its edges. We further demonstrate that this intervention yields more interpretable circuit structure. Using standard activation patching and circuit-discovery methods, we find that the sparsified models exhibit dramatically simplified computational graphs, reducing by up to two orders of magnitude the number of edges between components, while retaining essential algorithmic behaviors. Taken together, these results position attention sparsity as a practical inductive tool for surfacing the minimal functional backbone underlying model behavior.
Figure 1: Simple example showing the attention patterns (shown in blue) of sparse and non-sparse transformers trained on a two digit addition task. Both models are able to correctly predict the sum, but the attention patterns are very different: the non-sparse model solves the task with highly dispersed information flow, while the sparse model uses a highly interpretable attention pattern: in Layer 0, the model first attends to the corresponding digits to be added, then in Layer 1, it attends to the carry bit only if it is needed (see middle and right columns, where the model has to carry once and twice respectively).
2 Related Work
2.1 Sparse Attention
As self-attention is a key component of the ubiquitous Transformer architecture, a large number of variants of attention mechanisms have been explored in the literature. Related to our approach are sparse attention methods, which are primarily designed to alleviate the quadratic scaling of vanilla self-attention. These methods typically rely on masks based on fixed local and strided patterns [child2019generating] or sliding-window and global attention patterns [beltagy2020longformer, zaheer2020big] to constrain the receptive field of each token. While these approaches are successful in reducing the computational complexity of self-attention, they require hand-defined heuristics that do not reflect the internal computations learned by the model. Besides these fixed-pattern sparse attention methods, Top-$k$ attention, which enforces sparsity by dynamically selecting the $k$ most relevant keys per query based on their attention scores, has also been explored [gupta2021memory]. While Top-$k$ attention enables learnable sparse attention, the need to specify $k$ limits its scope for interpretability for two reasons. First, selecting the optimal $k$ is difficult, and setting $k$ too low can degrade model performance. Second, and more fundamentally, Top-k attention does not allow the model to choose different $k$ for different attention heads based on the context. We argue that this flexibility is crucial for maintaining model performance; indeed, we observe in our experiments (Sec. 2, for example) that our method induces attention heads that attend to a varying number of tokens and lead to emergent circuits that are sparser than even picking $k=1$, as some attention heads can be completely switched off. Instead of using Top-$k$ or hand-defined masks, we employ a sparsity regularisation approach, originally proposed in the context of world modelling [lei2025spartan], which serves as a soft inductive bias that flexibly induces sparse and interpretable circuits.
2.2 Mechanistic Interpretability
Mechanistic interpretability seeks to uncover how internal components of large language models (LLMs) implement specific computations. In Transformers, attention provides a natural handle for analysis, as attention maps reveal directed information flow between tokens and can often be linked to interpretable algorithmic roles [elhage2021mathematical]. Systematic ablation studies [nanda2023progress] provide a first step by measuring performance degradation when specific heads or neurons are removed. Activation patching tests how substituting activations from a source run into a target run affects model outputs [zhang2023towards]. Attribution patching [syed2024attribution] extends this framework by approximating patching effects via local linearisation, enabling scalable estimation of each component’s effective contribution. Together, these approaches allow researchers to isolate sub-circuits, minimal sets of attention heads and MLPs that are causally responsible for a given behavior or task [conmy2023towards]. Attention itself plays a dual role: it both routes information and exposes interpretable relational structure, making it a key substrate for mechanistic study. Our work builds on this foundation by leveraging sparsity to simplify these circuits, amplifying the interpretability of attention-mediated computation while preserving model performance.
While, to the best of our knowledge, our method is the first to demonstrate improved interpretability by regularising the sparsity of self-attention, several complementary works utilise sparsity as a driving principle to render other components of a Transformer interpretable. Focusing on the feed-forward MLPs, works such as Transcoders [dunefsky2024transcoders] and Cross-Layer Transcoders [ameisen2025circuit] have found success in uncovering interpretable structures by approximating MLPs in pretrained models using higher-dimensional MLPs with sparsely activated hidden neurons. Going beyond using sparsity as a post-hoc analysis tool, recent works have investigated sparsity regularisation during training to induce intrinsically interpretable models. [yang2025mixture], for example, replaces MLPs in the model architecture with mixtures of sparsely activated experts composed of interpretable neurons; Sparse Transformer weights [gao2024weightsparse] regularises model weights to enable the discovery of substantially smaller circuits.
3 Method
Our main hypothesis is that post-training existing LLMs to encourage sparse attention patterns leads to the emergence of more interpretable circuits. In order to instantiate this idea, we require a post-training pipeline that satisfies three main desiderata:
  1. 1.
    To induce sparse message passing between tokens, we need an attention mechanism that can ‘zero-out’ attention edges, which in turn enables effective L0-regularisation on the attention weights. This is in contrast to the standard softmax attention mechanism, where naive regularisation would result in small but non-zero attention weights that still allow information flow between tokens.
  2. 2.
    The model architecture needs to be compatible with the original LLM such that the pre-trained LLM weights can be directly loaded at initialisation.
  3. 3.
    The post-training procedure needs to ensure that the post-trained models do not lose prediction performance compared to their fully-connected counterparts.
To this end, we leverage the Sparse Transformer architecture in the SPARTAN framework proposed in [lei2025spartan], which uses sparsity-regularised hard attention instead of the standard softmax attention. In the following subsections, we describe the Sparse Transformer architecture and the optimisation setup, highlighting how this approach satisfies the above desiderata.
3.1 Sparse Attention Layer
Given a set of token embeddings, the Sparse Transformer layer computes the key, query, and value embeddings, $\{k_{i},q_{i},v_{i}\}$, via linear projections, analogous to the standard Transformer. Based on the embeddings, we sample a binary gating matrix from a learnable distribution parameterised by the keys and queries,
$A_{ij}\sim Bern(\sigma(q_{i}^{T}k_{j})),$ (1)
where $Bern(\cdot)$ is the Bernoulli distribution and $\sigma(\cdot)$ is the logistic sigmoid function. This sampling step can be made differentiable via the Gumbel Softmax trick [jang2017categorical]. This binary matrix acts as a mask that controls the information flow across tokens. Next, the message passing step is carried out in the same way as standard softmax attention, with the exception that we mask out the value embeddings using the sampled binary mask,
$SparseAttention(Q,K,V)=\bigg[A\odot softmax(\frac{QK^{T}}{\sqrt{d_{k}}})\bigg]V,$ (2)
where $d_{k}$ is the dimension of the key embeddings and $\odot$ denotes element-wise multiplication. During training, we regularise the expected number of edges between tokens based on the distribution over the gating matrix. Concretely, the expected number of edges for each layer can be calculated as
$\mathbb{E}\big[|A|\big]=\sum_{i,j}\sigma(q^{T}_{i}k_{j}).$ (3)
Note that during the forward pass, each entry of $A$ is a hard binary sample that zeros out attention edges, which serves as an effective L0 regularisation. Moreover, since the functional form of the sparse attention layer after the hard sampling step is the same as standard softmax attention, pre-trained model weights can be directly used without alterations. [^1] Technically, the sampled $A$ affects the computation. This can be mitigated by adding a bias term inside the sigmoid function to ensure all gates are open at initialisation. Experimentally, we found this to be unnecessary as the models quickly recover their original performance within a small number of gradient steps.
3.2 Constrained Optimisation
In order to ensure that the models do not lose prediction performance during the post-training procedure, as per desideratum 3, we follow the approach proposed in [lei2025spartan], which employs the GECO algorithm [rezende2018tamingvaes]. Originally developed in the context of regularising VAEs, the GECO algorithm places a constraint on the performance of the model and uses a Lagrangian multiplier to automatically find the right strength of regularisation during training. Concretely, we formulate the learning process as the following optimisation problem,
$\min_{\theta}\sum_{l}\mathbb{E}\big[|A_{l}|\big]\qquad s.t.\quad CE\leq\tau,$ (4)
where $A_{l}$ denotes the gating matrix at layer $l$, CECE is the standard next token prediction cross-entropy loss, and $\tau$ is the required target loss, and $\theta$ is the model parameters. In practice, we set this target as the loss of the pre-trained baseline models. We solve this optimisation problem via Lagrangian relaxation, yielding the following max-min objective,
$\max_{\lambda>0}\min_{\theta}\bigg[\sum_{l}\mathbb{E}\big[|A_{l}|\big]+\lambda(CE-\tau)\bigg].$ (5)
This can be solved by taking gradient steps on $\theta$ and $\lambda$ alternately. During training, updating $\lambda$ automatically balances the strength of the sparsity regularisation: when $CE$ is lower than the threshold, $\lambda$ decreases, and hence more weight is given to the sparsity regularisation term. This effectively acts as an adaptive schedule which continues to increase the strength of the regularisation until the model performance degrades.
4 Experiments
(a) Refer to caption
(b) Refer to caption
Figure 2: (a) Cross-entropy loss with respect to the mean proportion of non-zero attention edges across the evaluation dataset. (b) Example of attention patterns in GPT-2 compared with those in a sparse-attention GPT-2 model for a single IOI task instance.
To evaluate the effectiveness of our post-training pipeline, we finetune pre-trained LLMs and compare their prediction performance and interpretability before and after applying sparsity regularisation. In our experiments, two models, GPT-2 [radford2019language] (124M parameters) and LLaMA3-1B [grattafiori2024llama3herdmodels] (1B parameters), are finetuned on the OpenWebText dataset [Gokaslan2019OpenWeb]. Both models are finetuned on sequences of length 64, with the target cross-entropy set at 3.2 (for GPT-2) and 3.5 (for LLaMA), which are the cross-entropy loss achieved on the training set by the respective models. We focus on short context-windows in the current study for computational reasons. In the following evaluations, we aim to answer two key questions: (1) Can sparse models maintain comparable performance to their dense counterparts with sparse message-passing? (2) Do sparse models lead to smaller circuits that are more easily discoverable using standard mechanistic interpretability methods?
4.1 Sparse Attention
We first evaluate the performance and the achieved sparsity of the post-trained models. Figure 2 (left) shows the cross-entropy loss and the mean proportion of non-zero attention edges across the evaluation dataset. Here, both models are able to maintain comparable performance to their dense counterparts, with only a very small number of active attention edges: only 0.2% to 0.3% of the total edges are active on average. To further corroborate that the sparsity post-training effectively rewires the internal computation of LLMs in a non-trivial way, we also evaluate the dense baseline models using top-k attention, where, for each attention head, only the top-k attention weights are kept while the rest are set to zero. The results are shown in Figure 2 (left), which shows that the performance of dense models degrades significantly under top-k attention. This suggests that, in contrast to sparse models, the internal information flow of dense models is diffused across many attention edges. Remarkably, the average number of active edges in the sparse models is significantly lower than in the top-1 attention setting, indicating that some attention heads in the sparse models are not used at all. This is qualitatively demonstrated in Figure 2 (right), which shows an example attention activation for the sparse model, where a large number of attention heads are not attending to anything and hence do not contribute to the computation circuit. In summary, sparse post-training consolidates information flow into a small number of edges while maintaining the same level of performance. Next, we investigate whether these emergent sparse attention patterns yield smaller and more interpretable circuits under standard mechanistic interpretability techniques.
4.2 Interpretability
We focus on two classical circuit-discovery tasks: variants of the Indirect Object Identification (IOI) task and the Greater-Than task. For both tasks, we apply activation patching between correct and corrupted versions of each input. Activation patching yields an importance score for every attention head. These scores can be computed at two levels: (i) a single-sentence level, using a single pair of correct and corrupted inputs, and (ii) a global level, obtained by averaging scores across many task variants. In both settings, the resulting scores induce a ranking over attention heads based on their relative contribution to task performance. We then select the top-$k$ heads under each ranking and intervene on the model by freezing all remaining heads. In the standard dense-attention setting, we freeze non-selected heads to their mean activation, following prior work showing that this preserves model behaviour more faithfully [conmy2023towards]. In contrast, for the sparse-attention model we freeze non-selected heads to zero, which is both natural given the model’s inductive bias and simplifies the intervention procedure. We compare in Figure 6 the results between mean and zero ablation for completion.
(a) Refer to caption
(b) Refer to caption
Figure 3: (a) Logit attribution per sentence keeping only the top-$k$ attention heads selected individually for each sentence via activation patching scores. (b) Logit attribution using a fixed set of top-$k$ attention heads selected globally across all sentences by averaging activation patching scores.
For each value of $k$, we compute the logit difference between the correct and incorrect answers under the intervention, and compare it to the clean model’s logit difference. We analyse the difference of logit differences and normalise the resulting curves by dividing by the absolute value of the first point and inverting the scale, yielding values between 0 and 1. We present normalised curves for both single-sentence top-$k$ head rankings and global top-$k$ rankings. The single-sentence analysis captures the number of heads needed to preserve performance on a specific prompt, while the global analysis reflects the stability and consistency of the underlying circuit across many task instances.
[…]
(a) Refer to caption       (b) Refer to caption
Figure 4: (a) Cumulative distribution of the sorted attribution-patching edge scores. The dotted lines show the cumulative distribution of the mean scores across sentences, reflecting circuit stability. The solid lines show the mean cumulative distribution of the scores across sentences. The plots for LLaMA are included in Appendix B (b) Example of the attention-head edges required to reach 0.9 cumulative score based on the averaged scores for the IOI task, representing the general circuit across all sentences. The comparison is shown for GPT-2 and its sparse-attention variant.
[…]
5 Discussion
Overall, our results indicate that applying sparsity to attention post-training can significantly reduce the number of active attention edges while keeping the model’s original loss unchanged. Effectively, the proposed post-training paradigm serves as a "cleaning" step for pre-trained large models, where the information flow through the model is reorganised into smaller and modular circuits that are much easier to interpret and analyse for downstream mechanistic interpretability techniques. Our experiments show that sparsity makes the model more interpretable at the circuit level by reducing the number of components involved in particular task behaviours.
The strength of our approach lies in its flexibility: the sparse attention layer serves as a direct slot-in replacement in any Transformer-based model, while the constrained optimisation scheme can be applied to any arbitrary loss function. As such, our post-training strategy naturally lends itself to a plethora of exciting potential extensions for future research. First, more advanced mechanistic interpretability techniques can be explored. In particular, we believe that sparse attention can play an important role in enabling attention attribution in cross-layer analyses and in circuit-tracing approaches  [kamath2025tracing, hanna2025circuit]. Another potential extension of our work is to explore other fine-tuning techniques such as low-rank finetuning or distillation, in order to reduce the computational and data requirements [hu2022lora]. Finally, while our experiments have demonstrated the efficacy of sparsity regularisation on models of up to 1B parameters, future work should investigate whether our findings extend to frontier models, potentially with the help of the aforementioned distillation and LoRA techniques.
6 Conclusion
Achieving interpretability requires innovations in both interpretation techniques and model design. In this work, we investigate how large models can be trained to be intrinsically interpretable. We present a flexible post-training procedure that sparsifies transformer attention while preserving the original pretraining loss. By minimally adapting the architecture, we apply a sparsity penalty under a constrained-loss objective, allowing pre-trained model to reorganise its connectivity into a much more selective and structured pattern.
Mechanistically, this sparsity translates into far simpler circuits: task-relevant computation concentrates into a small, stable set of heads and edges. On two typical tasks, IOI and greater-than, we can recover most of the models’ behaviour from circuits that are orders of magnitude smaller than in the dense case. This positions post-hoc sparse attention as a practical tool for “cleaning up” pretrained models, and suggests promising directions for extending sparsity-based interventions to other components and tasks to further enhance mechanistic interpretability.
Appendix A Task Definitions: Greater-Than and IOI
A.1 Greater-Than Task
Each example contains a clean prompt, a corrupt prompt, and two disjoint sets of candidate continuations, answers and wrong_answers. A typical entry is:
  "clean":   "The demonstrations lasted from the year 1363 to 13",
  "corrupt": "The demonstrations lasted from the year 1301 to 13",
  "answers": ["64", "65", ..., "99"],
  "wrong_answers": ["00", "01", ..., "63"]
For the clean prompt, any token in answers yields an end year strictly greater than the start year (e.g. "1364""1399"), whereas tokens in wrong_answers correspond to years that are less than or equal to the start year. The corrupt prompt changes only the starting year, shifting which continuations correspond to valid end years. We use the logit difference between the aggregated probability mass on answers vs. wrong_answers in clean vs. corrupt contexts as our signal, in the spirit of prior mechanistic studies on simple algorithmic tasks [elhage2021mathematical, nanda2023progress]. For LLaMA-1B, the tokenizer merges the first three digits of the year into a single token, which prevents the model from solving the task. We therefore simplify the task by requiring the model to predict only the final digit of the end year.
A.2 Indirect Object Identification (IOI)
Our IOI setup follows the standard indirect object identification paradigm for mechanistic interpretability [elhage2021mathematical, conmy2023towards]. Each example is generated by combining:
  • a pair of names $(A,B)$, e.g. (" Mary", " John");
  • a natural-language template with placeholders [A], [B], and [S].
We instantiate templates such as:
"Then, [B] and [A] went to the park. [S] gave a ball to"
"When [B] and [A] got a snack at the cafe, [S] decided to give it to"
"After the lunch, [B] and [A] went to the mall. [S] gave a gift to"
by sampling a name pair and substituting $[A]$ and $[B]$, then choosing the subject $[S]$ (either one of the pair). The correct continuation is the indirect object, i.e. the other member of the pair.
For example, with $(A,B)=(\texttt{" John"},\texttt{" Mary"})$ and $S=B$, one instance is:
Then, Mary and John went to the park. Mary gave a ball to
The correct continuation is " John", while " Mary" and any distractor names are treated as incorrect candidates.
Appendix B Additional results
Figure 5: Cumulative distribution of the sorted attribution-patching edge scores. The dotted lines show the cumulative distribution of the mean scores across sentences, reflecting circuit stability. The solid lines show the mean cumulative distribution of the scores across sentences.
Figure 6: Sentence-level logit attribution using top-$k$ components selected per sentence via activation patching, comparing mean-activation ablation against zero-activation ablation.